OpenStudy (anonymous):

how to get inverse laplace transform of this equation s^2+ squrt 2s + 1

OpenStudy (anonymous):

is it... \[s ^{2} +\sqrt{2s} + 1\]

OpenStudy (anonymous):

yes the whole equation in denominator

OpenStudy (anonymous):

\[\sqrt{2} e^{-t/\sqrt{2}} \sin (t/\sqrt{2})\] this is ans but i want explanation of middle steps shortly

OpenStudy (anonymous):

i'm still working on simplifying the equation... :)

OpenStudy (anonymous):

isn't it... \[\frac{ 1 }{ s ^{2} + \sqrt{2} s + 1 } ?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i think i hav to remember all laplace transform formulae Thanks for u r effort what s u r profession ....?

OpenStudy (anonymous):

i'm an electronics engineer by profession...

OpenStudy (anonymous):

sorry i was not able to finish your problem... got busy this day in my work... by the way... it's really the table such as... \[L ^{-1}\left\{ \frac{ a }{ (s ^{2}+a ^{2}) } \right\}=\sin at\]

OpenStudy (anonymous):

another is... \[L ^{-1}\left\{ \frac{ 1 }{ (s+a) } \right\}=e ^{-at}\]